# On Long-Term Problems in Multiplicative Ideal Theory and Factorization Theory

### v2

@@ 12.1. Half-Factoriality
In 1960 Carlitz proved that a ring of integers in an algebraic number field is half-factorial if and only if its class group has at most two elements. Ever since then half-factoriality is a key property studied in factorization theory and an abundance of results has been achieved (we refer to the surveys [29, 34] for some early work). We can pick only one recent highlight and one of our favorite problems.
@@ 12.2. Finiteness of Arithmetic Invariants
The results of (1) can be found in [64] and (2) follows from [62, Theorem 4.2]. ∎
@@
$M$-equivalence defines a congruence relation on $F$ and for $y\in F$, let $[y]_{M}^{F}=[y]$ denote the congruence class of $y$. Then
@@
    $M$ is Mori, $(M\negthinspace:\negthinspace\widehat{M})\neq\emptyset$, and $\mathcal{C}_{v}(\widehat{M})$ is finite.
@@
Let $M$ be a Mori monoid with $(M\negthinspace:\negthinspace\widehat{M})\neq\emptyset$ such that $\mathcal{C}_{v}(\widehat{M})$ is finite. Characterize when $M$ is a C-monoid. The cases when $M$ is seminormal or strong Mori are of special interest ([74]).
@@
    Let $M$ be a monoid such that $M_{{\text{\rm red}}}$ is finitely generated. Then $M$ is Mori with $(M\negthinspace:\negthinspace\widehat{M})\neq\emptyset$, and $M$ is a C-monoid if and only if $\mathcal{C}_{v}(\widehat{M})$ is finite.
@@
    Let $D$ be Mori domain with $(D\negthinspace:\negthinspace\widehat{D})\neq\{0\}$ and with $\mathcal{C}_{v}(\widehat{D})$ finite. If $\widehat{D}/(D\negthinspace:\negthinspace\widehat{D})$ is finite, then $D$ is a C-domain. The converse holds if $D$ is semilocal, Noetherian but not local.
@@
For Part (1), we refer to [66, Proposition 4.8]. The first statement in Part (2) follows from [64] and the converse statement is proved by Reinhart in [127, Corollary 4.5]. ∎
Every generalized Cohen-Kaplansky domain (this is an atomic domain that has only finitely many atoms, up to associates, that are not prime) is a C-domain ([71, Proposition 4.11]). Part 2 of Theorem 12.4 shows that, in particular, orders in algebraic number fields are C-domains. A classical result of Furtwängler characterizes the ideals that can be realized as conductor ideals of an order ([87, Theorem 2.12.10]). This was generalized by Reinhart to a more general setting of Noetherian domains ([128]), and his result gives rise to the following problem.
@@
Let $H$ be a Krull monoid with finite $v$-class group. Characterize the ideals $\mathfrak{f}\subseteq H$ for which there is a C-monoid $M\subseteq H$ with $\widehat{M}=H$ and with conductor $\mathfrak{f}=(M\negthinspace:\negthinspace\widehat{M})$.
@@ 12.3. The Transfer Krull Property
Along the way, it was observed that there are monoids and domains, that are not Krull, but nevertheless they allow transfer homomorphisms to monoids of zero-sum sequences. Inspired by the success of the powerful methods from additive combinatorics, intense investigations have been started to figure out under which conditions we do have transfer homomorphisms to monoids of zero-sum sequences. Deep results revealed that this may happen even for non-commutative rings ([136]). Here we restrict our discussion to monoids (so, to commutative cancellative semigroups) and just mention two results. A one-dimensional local Mori domain $D$ is transfer Krull if and only if $D$ is half-factorial (whence in the trivial case only; see [73, Theorem 5.5]). Recent characterizations for orders in Dedekind domains are given in [125]. For a survey and more, see [15, 75, 14].
@@
Let $R$ be a not necessarily commutative ring and let $\mathcal{C}$ denote a class of left $R$-modules, that is closed under direct sums, direct summands, and isomorphisms. Let $\mathcal{V}(\mathcal{C})$ be the semigroup of isomorphism classes of modules from $\mathcal{C}$, with operation induced by the direct sum. If every module from $\mathcal{C}$ has a semilocal endomorphism ring, then $\mathcal{V}(\mathcal{C})$ is Krull (this was proved by Facchini; for examples, see [44]). For a given class $\mathcal{C}$, for which $\mathcal{V}(\mathcal{C})$ is Krull, study classes $\mathcal{C}^{\prime}$ of modules having a transfer homomorphism $\theta\colon\mathcal{V}(\mathcal{C}^{\prime})\to\mathcal{V}(\mathcal{C})$. For a simple (artificial) example, of how this strategy could work, see [75, Example 5.4].
@@ 12.4. What can a family of length sets look like?
In contrast to Theorem 12.2 there is the following result by Kainrath (see [64, Theorem 7.4.1]).
@@
The system $\mathcal{L}(M)$ gathers all length sets of a given monoid $M$ and Theorems 12.2 and 12.5 describe all the phenomena that might occur. But, in addition, one might ask after the typical structure of a length set when we pick a random element. Let $M$ be an analytic monoid. This is a Krull monoid with finite $v$-class group and with a norm function that allows to establish a theory of L-functions ([98]). Then almost all (in the sense of Dirichlet density) length sets are intervals ([64, Theorem 9.4.11]). Other concepts of density include the Zariski density (say for $\mathbb{Q}$-algebras). There is much heuristic evidence (in a variety of settings) that most (in what sense ever) length sets are much simpler compared with the possible structure of all length sets popping up in $\mathcal{L}(M)$. Thus, we present the following problem.
@@
We end with a short discussion on semigroups of ideals that are not necessarily cancellative. The algebraic and arithmetic structure of ideal semigroups has found wide attention in the literature (for a sample of recent contributions, see [96, 3, 48, 71, 18, 97, 69, 58, 108]). In general, these semigroups need not be cancellative (indeed, the semigroup of all nonzero ideals of a domain is cancellative if and only if the domain is almost Dedekind [84, Theorem 23.2]). However, semigroups of ideals are unit-cancellative in a variety of significant cases. To discuss this situation, let $M$ be a commutative semigroup with identity. We say that $M$ is unit-cancellative if $a,u\in M$ and $au=a$ implies that $u\in H^{\times}$. Factorization theory in unit-cancellative monoids can be developed along the same lines as in the cancellative setting. In particular, atoms and length sets are defined as it is done for monoids in Section 2.

### v3

@@ 12.1. Half-Factoriality
In 1960 Carlitz proved that a ring of integers in an algebraic number field is half-factorial if and only if its class group has at most two elements. Ever since then half-factoriality is a key property studied in factorization theory and an abundance of results has been achieved (we refer to the surveys [29, 34] for some early work and to [38] for a recent characterization of half-factoriality). We can pick only one recent highlight and one of our favorite problems.
@@ 12.2. Finiteness of Arithmetic Invariants
The results of (1) can be found in [66] and (2) follows from [64, Theorem 4.2]. ∎
@@
$M$-equivalence defines a congruence relation on $F$ and for $y\in F$, let $y]_{M}^{F}=[y]$ denote the congruence class of $y$. Then
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
For Part (1), we refer to [68, Proposition 4.8]. The first statement in Part (2) follows from [66] and the converse statement is proved by Reinhart in [128, Corollary 4.5]. ∎
Every generalized Cohen-Kaplansky domain (this is an atomic domain that has only finitely many atoms, up to associates, that are not prime) is a C-domain ([73, Proposition 4.11]). Part 2 of Theorem 12.4 shows that, in particular, orders in algebraic number fields are C-domains. A classical result of Furtwängler characterizes the ideals that can be realized as conductor ideals of an order ([89, Theorem 2.12.10]). This was generalized by Reinhart to a more general setting of Noetherian domains ([129]), and his result gives rise to the following problem.
@@
[…]
@@ 12.3. The Transfer Krull Property
Along the way, it was observed that there are monoids and domains, that are not Krull, but nevertheless they allow transfer homomorphisms to monoids of zero-sum sequences. Inspired by the success of the powerful methods from additive combinatorics, intense investigations have been started to figure out under which conditions we do have transfer homomorphisms to monoids of zero-sum sequences. Deep results revealed that this may happen even for non-commutative rings ([137]). Here we restrict our discussion to monoids (so, to commutative cancellative semigroups) and just mention two results. A one-dimensional local Mori domain $D$ is transfer Krull if and only if $D$ is half-factorial (whence in the trivial case only; see [75, Theorem 5.5]). Recent characterizations for orders in Dedekind domains are given in [126]. For a survey and more, see [15, 77, 14].
@@
Let $R$ be a not necessarily commutative ring and let $\mathcal{C}$ denote a class of left $R$-modules, that is closed under direct sums, direct summands, and isomorphisms. Let $\mathcal{V}(\mathcal{C})$ be the semigroup of isomorphism classes of modules from $\mathcal{C}$, with operation induced by the direct sum. If every module from $\mathcal{C}$ has a semilocal endomorphism ring, then $\mathcal{V}(\mathcal{C})$ is Krull (this was proved by Facchini; for examples, see [46]). For a given class $\mathcal{C}$, for which $\mathcal{V}(\mathcal{C})$ is Krull, study classes $\mathcal{C}^{\prime}$ of modules having a transfer homomorphism $\theta\colon\mathcal{V}(\mathcal{C}^{\prime})\to\mathcal{V}(\mathcal{C})$. For a simple (artificial) example, of how this strategy could work, see [77, Example 5.4].
@@ 12.4. What can a family of length sets look like?
In contrast to Theorem 12.2 there is the following result by Kainrath (see [66, Theorem 7.4.1]).
@@
The system $\mathcal{L}(M)$ gathers all length sets of a given monoid $M$ and Theorems 12.2 and 12.5 describe all the phenomena that might occur. But, in addition, one might ask after the typical structure of a length set when we pick a random element. Let $M$ be an analytic monoid. This is a Krull monoid with finite $v$-class group and with a norm function that allows to establish a theory of L-functions ([100]). Then almost all (in the sense of Dirichlet density) length sets are intervals ([66, Theorem 9.4.11]). Other concepts of density include the Zariski density (say for $\mathbb{Q}$-algebras). There is much heuristic evidence (in a variety of settings) that most (in what sense ever) length sets are much simpler compared with the possible structure of all length sets popping up in $\mathcal{L}(M)$. Thus, we present the following problem.
@@
We end with a short discussion on semigroups of ideals that are not necessarily cancellative. The algebraic and arithmetic structure of ideal semigroups has found wide attention in the literature (for a sample of recent contributions, see [98, 3, 50, 73, 18, 99, 71, 60, 110]). In general, these semigroups need not be cancellative (indeed, the semigroup of all nonzero ideals of a domain is cancellative if and only if the domain is almost Dedekind [86, Theorem 23.2]). However, semigroups of ideals are unit-cancellative in a variety of significant cases. To discuss this situation, let $M$ be a commutative semigroup with identity. We say that $M$ is unit-cancellative if $a,u\in M$ and $au=a$ implies that $u\in H^{\times}$. Factorization theory in unit-cancellative monoids can be developed along the same lines as in the cancellative setting. In particular, atoms and length sets are defined as it is done for monoids in Section 2.